OpenStudy (anonymous):

How do I solve this question? 8x-9=15

OpenStudy (anonymous):

8x = 15+9, so 8x = 24, x = 3

OpenStudy (anonymous):

Can you please break that up some more?

OpenStudy (anonymous):

ok. 8x - 9 = 15. so 8x - 9 + 9 = 15 + 9 (adding 9 to both sides). 15 + 9 = 24 and -9 + 9 = 0. So 8x = 24. Dividing both sides by 8, you get 8x / 8 = 24 / 8 , and so 8x / 8 = x and 24/8 = 3. So x = 3.

OpenStudy (anonymous):

Thanks for the help!!!